OpenStudy (anonymous):

Pb(NO3)2(aq) + 2KI(aq) → PbI2(s) + 2KNO3(aq 3 CuSO4 + 2 Al → Al2(SO4)3 + 3 Cu MgO + CO2 → MgCO3 2 H2O(l) → 2 H2(g) + O2(g) MgCl2 + Ca(OH)2 → Mg(OH)2 + CaCl2 One of the reactions you conducted can actually be categorized as TWO different types of reactions. Which reaction is this, and what are the two types of reactions?

OpenStudy (kira_yamato):

Pb(NO3)2(aq) + 2KI(aq) → PbI2(s) + 2KNO3(aq) = Precipitation rxn 3 CuSO4 + 2 Al → Al2(SO4)3 + 3 Cu = Displacement rxn / Redox MgO + CO2 → MgCO3 = Acid-Base (sort of) 2 H2O(l) → 2 H2(g) + O2(g) = Decomposition / Disproportionation MgCl2 + Ca(OH)2 → Mg(OH)2 + CaCl2 = Precipitation rxn